OpenStudy (anonymous):

Find the exact solution log5(x-4)=-1. I got an answer of -1 is this right?

OpenStudy (lalaly):

\[\log_5(x-4)=-1\]\[5^{-1}=x-4\]

OpenStudy (lalaly):

\[x=\frac{1}{5}+4\]

OpenStudy (anonymous):

where did the 1/5 come form and why is it not negative

OpenStudy (lalaly):

\[5^{-1}=\frac{1}{5}\]

OpenStudy (anonymous):

oh

OpenStudy (lalaly):

\[\large{x^{-n}=\frac{1}{x^n}}\]

OpenStudy (anonymous):

so the end answer is 4.2?

OpenStudy (lalaly):

yes

OpenStudy (anonymous):

awesome i got it so hold on thers one more

OpenStudy (anonymous):

if the base of log is 10 then answer will be 4.02 and if the base is 5 then answer will be 4.2